Chapter 9/10/11 , 2019 Based upon the following facts calculate the Weighted Average Cost of Capital (WACC) for Student Success Corporation (SSC): PART 1 – WACC  Tax rate = 40%  Debt Financing: $10,000 Face Value 10-Year, 5% Coupon, Semiannual Non-Callable Bonds Selling for $11,040 New bonds will be privately placed with no flotation cost.  Common Stock: Current Price $40; Current Dividend = $3.00 and Growth Rate = 5%.  Common Stock: Beta = 1.1; Risk Free Rate 2.0%; Required Return of the Market 7%  Capital structure: 40% Debt, 60% Common Equity 1) What is the cost of debt? 2) What is the cost of equity – use both CAPM and the Dividend Model? 3) What is the WACC – for Equity you can use either answer above or an average? PART 2 – Capital Budget If SSC is deciding upon whether to approve a capital project and the cash flows are as follows: Year 0 (initial investment) $2,000 Year 1 Cash Flow $1,000 Year 2 Cash Flow $600 Year 3 Cash Flow $400 Year 4 Cash Flow $4,000 Calculate: 1) Net Present Value (NPV) ____________________________ 2) Internal Rate of Return (IRR) ____________________________ 3) Payback Period ____________________________ Remember to use the WACC from above for the NPV and to compare to the IRR. Should this project be approved and why or why not?
Answer #1

1. Cost of Debt before tax calculated using Yield To Maturity Approach :-

1-(1+r) BondPrice = Semiannual Interest * - FaceValue +- + (1+r)

10000 11040 = 250 * ,1- (1+r) -20 - ) + (1 + r20

By solving the above equation, we get r = 1.8718%. Thus, Before Tax Cost of Debt is 3.7436% (1.8718*2)

After Tax Cost of Debt = 3.7436% * (1-0.4) = 2.2462%

2. Cost of Equity using CAPM Model = Risk Free Rate + Beta* Risk Premium

                                                           = 2 + 1.1* (7-2)

                                                            = 7.5%

Cost of Equity using Dividend Model = (Dividend/ Current Market Price) +Growth Rate

                                                          = (3/40) + 5%

                                                         = 12.5%

Cost of Equity using average of above two is 10% ((7.5+12.5)/2).

3. WACC if 40% Debt and 60% Equity = (2.2462*0.4)+(10*0.6) = 6.8985%.

Part 2

1. NPV Calculation :-

Year Cash Flow Discounting Factor @ 6.8985% Present Value
0    (2,000.00)                       1.0000        (2,000.00)
1      1,000.00                       0.9355              935.47
2          600.00                       0.8751              525.06
3          400.00                       0.8186              327.45
4      4,000.00                       0.7658          3,063.19
Total          2,851.16

2. IRR Calculation :-

Year Cash Flow Discounting Factor @ 45% Present Value Discounting Factor @ 46% Present Value
0    (2,000.00)                       1.0000        (2,000.00) 1           (2,000.00)
1      1,000.00                       0.6897              689.66 0.684931507                 684.93
2          600.00                       0.4756              285.37 0.469131169                 281.48
3          400.00                       0.3280              131.21 0.321322719                 128.53
4      4,000.00                       0.2262              904.87 0.220084054                 880.34
Total                11.11                 (24.72)

By solving the above two rate, we get IRR as 45.31%.

3. Payback Period

Year Cash Flow Cumulative Cash Flows
0    (2,000.00)                 (2,000.00)
1      1,000.00                 (1,000.00)
2          600.00                     (400.00)
3          400.00                                 -  
4      4,000.00                    4,000.00

Thus, the payback period is 3 years for the project.